1. Amna wants to purchase a used Lexus for OR 20,000.00. She could pay OR 5000.00 at purchase and wishes to pay the balance in 5 yearly equal installments at a 12% interest. How much would Amna’s yearly installments payment be?

2. Mohammed wants to buy a house in another seven years. The type of house he wants to buy is currently selling for OR 50,000, and the price will increase at a compound rate of 10% per year. Mohammed can presently invest in a bank fixed deposit earning a compound annual rate 14% per year. How much money must Mohammed deposit in the bank each year to be able to purchase a house in seven years from now?

3. Covid-19 virus has created health uncertainties. You could buy an health insurance for yourself against potential Covid viral infection by paying OR 24 every year or, you may buy a lifetime insurance for a onetime payment of OR 400 today. Using a discount rate of 5%, how many years does it take to make the lifetime subscription the better option compared to yearly payments?
Answer #1

1.
=(20000-5000)*12%/(1-1/1.12^5)=4161.14597911573

2.
=50000*1.1^7*14%/(1.14^7-1)=9080.27896147222

3.
=NPER(5%,24,-400)=36.7237843883015